OpenStudy (castiel):

Why does integral of sin(sqrtx)dx has to be done with integration by parts and not just substitution or advanced guessing?

OpenStudy (anonymous):

\[u=\sqrt x~~\Rightarrow~~du=\frac{1}{2\sqrt x}~dx~~\iff~~2u~du=dx\] So your integral can be written as \[\int\sin\sqrt x~dx=2\int u\sin u~du\] the latter which requires integration by parts.